OpenStudy (anonymous):

Explain why the sum, the difference, and the product of two rational numbers are again rational numbers. Is the product of two irrational numbers again irrational? What about the sum?

OpenStudy (anonymous):

please somebody help me :(

OpenStudy (perl):

hello

OpenStudy (perl):

the sum of two rational numbers is rational , proof: given p/q, r/s , q and s are both non zero (because division by zero is undefined). now is p/q + r/s = a rational # ?

OpenStudy (perl):

yes it is , because p/q + r/s = ( ps + rq) / (q*s), and we assumed both q and s are non-zero. and ps+rq is an integer, qs is an integer. so this is a ratio of two integers, so this is rational

OpenStudy (perl):

the sum and product of two irrational numbers is not necessarily irrational (3+sqrt(2) + (3-sqrt(2)) = 6 sqrt(2) * sqrt(2) = 2

OpenStudy (anonymous):

how about the difference and the product?

OpenStudy (anonymous):

thanks alot :)

OpenStudy (perl):

sure thing

OpenStudy (anonymous):

ummm..sorry..can you explain why?

OpenStudy (anonymous):

hello?

OpenStudy (anonymous):

product of two irrational numbers is a rational number some times like

OpenStudy (anonymous):